satellite73 (satellite73):

\[\lim_{x \to\infty}xe^{\frac{ 1 }{ x }}-x\]

OpenStudy (anonymous):

or more to the point, why is it 1?

OpenStudy (anonymous):

holy cow! wut is this math

OpenStudy (anonymous):

or more more to the point, why is it 1 without using power series?

OpenStudy (anonymous):

Hum did you try finding the first derivative?

jimthompson5910 (jim_thompson5910):

\[\Large \lim_{x\to \infty}(xe^{1/x} - x)\] \[\Large \lim_{x\to \infty}(x(e^{1/x} - 1))\] \[\Large \lim_{x\to \infty}\left(\frac{e^{1/x} - 1}{1/x}\right)\] then use L'Hospitals rule

OpenStudy (anonymous):

oh wow thanks !

jimthompson5910 (jim_thompson5910):

no problem

OpenStudy (anonymous):

old trick i forgot to use here the form was messing me up

jimthompson5910 (jim_thompson5910):

yeah I had to look it up myself a while back. It has to be in the form f/g

OpenStudy (tylerd):

thanks guys